Mathematics 241 – Multivariable Calculus Quiz 8 – November 19, 2010

Your Name:

Let R = {(x, y) ∈ R^{2} : y ≥ |x| and y ≤ 3} and let f(x, y) be a general continuous function on R.

A) (5) Sketch the region R.

B) (7.5) Find the limits of integration for iterated integral(s) of the form ∫ ∫ f(x, y) dy dx to compute ∫ ∫_{R} f(x, y) dA.

C) (7.5) Compute ∫ ∫_{R} x dA.